OpenStudy (anonymous):

Given that: \[\int\limits_{0}^{3} 1/3 f(x) dx = 5 \] find the value of: \[\int\limits_{3}^{0} f(x) dx \]

OpenStudy (ikram002p):

\(\color{blue}{\Huge \int\limits_{0}^{3} \frac{1}{3} f(x) dx = 5}\) \(\large \int\limits_{0}^{3} \frac{1}{3} f(x) dx =- \int\limits_{3}^{0} \frac{1}{3} f(x) dx\) =\(\large-\frac{1}{3} \int\limits_{3}^{0} f(x) dx\)

OpenStudy (ikram002p):

\(\large -\frac{1}{3}\int\limits_{0}^{3} f(x) dx = 5 \) \(\large \int\limits_{0}^{3} f(x) dx = -5×3=-15 \) got it ?

OpenStudy (anonymous):

understand. but where the negative 1/3 come from? is it bcz we turn it upside down so automatically it become negative integer

OpenStudy (ikram002p):

yess ur correct !

OpenStudy (anonymous):

wah, tq tq tq. u help me alot.

OpenStudy (ikram002p):

np, ur wlc ^^